OpenStudy (anonymous):

Which of the following is the explicit formula for the sequence below? 8, 12, 16, 20, ... A. an = 8 + (n - 1)4 B. an = 8 + 4n C. an = 4 + (n - 1)8 D. an = 4 + 8n

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

start by letting n=1; which ones equal 8 as a result?

OpenStudy (anonymous):

A and B ?

OpenStudy (anonymous):

would it be A ?

hartnn (hartnn):

only A results in 1st term =8, so yes, its A

OpenStudy (anonymous):

thank you @hartnn

OpenStudy (amistre64):

B would equal 12 :)

OpenStudy (amistre64):

do you know what your text suggests you start your ns with? 0 or 1?

OpenStudy (amistre64):

8 + 4n 8 + 4.0 = 8 8 + 4.1 = 12 8 + 4.2 = 16 8 + 4.3 = 20 or .... 8 + (n - 1)4 8 + (1 - 1)4 = 8 8 + (2 - 1)4 = 12 8 + (3 - 1)4 = 16 8 + (4 - 1)4 = 20

OpenStudy (amistre64):

so this all depends on what your material defines for the starting point of "n"

OpenStudy (anonymous):

thank you ! :) that helps alot !

OpenStudy (amistre64):

good luck :)